### v2

# Counting observables in stochastic excursions

Abstract
Understanding fluctuations of observables across stochastic trajectories is essential for various fields of research, from quantum thermal machines to biological motors. We introduce the notion of stochastic excursions as a framework to analyze sub-trajectories of processes far from equilibrium. Given a partition of state space in two phases, labeled active and inactive, an excursion starts with a transition into the active phase and ends upon returning to inactivity. By incorporating counting observables, our approach captures finite-time fluctuations and trajectory-level behavior, providing insights on thermodynamic trade-offs between energy expenditure, entropy production and dynamical activity. As our main result, we uncover a fundamental relation between fluctuations of counting observables at the single-excursion level and the steady state noise obtained from full counting statistics. We also show the existence of an exchange-type fluctuation theorem at the level of individual excursions. As an application, we explore how analyzing excursions yields additional insights into the operation of the three-qubit absorption refrigerator.
Introduction—A wide variety of phenomena in nature can be characterized by two alternating phases: an “inactive” phase $A$ and an “active” phase $B$. Every once in a while the system becomes activated (e.g., when some energy enters the system), transitioning from $A\to B$. It then spends some random time in $B$ (e.g. performing some task) and eventually returns from $B\to A$, see Fig. 1(a). We term this sequence of events a stochastic excursion Chung1976 ; Majumdar2005 ; Majumdar2008 ; Majumdar2015 ; Takacs1991 ; Mazzolo2017 . Problems of this form have been widely studied across various fields of science, including biology bergPhysicsChemoreception1977 ; Mora2010 ; Fancher2020 ; Harvey2023 ; Barato2015 ; Lang2014 , physics Prech2024 ; Hegde2025 , chemistry bialekPhysicalLimitsBiochemical2005 , and applied mathematics kleinrock1974queueing ; Lindley1952 ; gross2011fundamentals ; Bhat2008 . Most theoretical frameworks to date are model-oriented and have focused on questions related to excursion times; e.g., how long, on average, will an excursion last, or what is the distribution of inactive periods between two active periods. These are all generalizations of first-passage time problems Yao1985 ; Rubino1989 ; Kook1993 ; Takacs1974 ; Garrahan2017 ; Bebon2024 .
Figure 1: (a) A system undergoes stochastic dynamics on a finite set of states. The states are split into two regions, called $A$ (gray) and $B$ (green). An excursion starts with a transition from $A\to B$ and ends with the first transition back from $B\to A$ (red arrows). The states $A$ do not have to be necessarily connected and clustered, they can be spread out across the state space. (b) Energy landscape of the three qubit refrigerator model, where the arrows indicate the possible transitions. Colors indicate which reservoir induced the transition. Region $A$ consists of the ground state (gray) and $B$ of all excited states (colored). We adopt the convention “$chw$” to denote the states. (c) Diagram of the three qubit absorption refrigerator, each qubit is connected to its respective heat reservoir and has a different energy gap. Labels are $c$ for cold, $h$ for hot, and $w$ for work. The resonant interaction $g$ couples the states $|010\rangle\leftrightarrow|101\rangle$ and takes place only if $\omega_{c}+\omega_{w}=\omega_{h}$. (d) The statistics of the excursion duration and counting variables, irrespective of the exact trajectory within $B$, can be viewed as a problem of inferring properties of $B$ from the empirically available $\{\hat{\mathcal{T}},\hat{Q}\}$. (e) Cellular sensing is often modelled as a stochastic process where some states represent unbound receptors while others represent bound, properties such as the environmental concentration of ligands are learned from the excursions throughout signaling states.
Our goal in this Letter is to provide a model-agnostic framework to analyze processes far from equilibrium, by focusing on the statistics of counting observables during excursions and their interplay with excursion times. These can be used to quantify and characterize any figure of merit that is linear in the number of transitions, such as thermodynamic currents, energy expenditure, entropy production, dynamical activity, and work extraction. As we show, this will allow us to address current fluctuations at the level of individual excursions, providing novel insights on the mechanisms responsible for the noise in any continuous-time Markov process. In particular, we unveil a new relation between fluctuations at the level of a single excursion and the steady-state noise obtained from full counting statistics (FCS) Esposito2007 ; Esposito2009 ; Brandes2008 ; Saito2008 ; landi2024a . We also show how individual excursions satisfy an exchange fluctuation theorem Jarzynski1997 ; Crooks1999 ; Talkner2007a ; Talkner2007b ; Talkner2008 ; Hasegawa2019a ; Hasegawa2019b ; Campisi2009 ; Campisi2010 ; Campisi2014 ; Andrieux2009 ; Jarzynski2004 ; Jarzynski2011 ; Seifert2012 ; Esposito2009 ; Harris2007 ; Horowitz2019 ; Nakamura2010 . In this Letter, the focus is to introduce the framework and discuss overarching ideas related to stochastic excursions. In an accompanying paper supp , we provide a suite of technical results on how to efficiently compute excursion-related quantities, and also discuss various relevant examples.
Prevalence of excursions—Specializing counting observables to stochastic excursions is relevant due to the ubiquitous role of excursions across numerous problems of interest. Below is a non-exhaustive list of examples spanning various disciplines where excursions are tightly connected to the main questions.
Absorption refrigerators. Refs. aamirThermallyDrivenQuantum2025 ; blokQuantumThermodynamicsQuantum2025 recently provided an experimental demonstration of a quantum absorption refrigerator. The system consists of three qubits each connected to their respective reservoirs, denoted $c$ (cold), $h$ (hot), and $w$ (work), see Fig. 1(c). The desired task is to cool down the cold qubit. The three qubits interact resonantly through an effective three-body interaction, allowing the conversion of an excitation in the work and cold reservoirs to one in the hot reservoir, $1_{c}0_{h}1_{w}\leftrightarrow 0_{c}1_{h}0_{w}$, effectively cooling the cold qubit. Region $A$ could represent the subset of states where the three qubits are in the ground state. $B$ represents everything else, as in Fig. 1(b). By constructing counting observables over excursions in $B$, one can fully explore the energetics of this model. We can unravel the refrigerator operation into excursions that represent successes and failures, whose statistics shed light on the overall performance. Moreover, we can characterize the fluctuations by decomposing the noise in terms of statistical contributions of the individual excursions. We illustrate this result by studying fluctuations of entropy production and dynamical activity. In this Letter, we discuss a simplified version of the absorption refrigerator model, while a detailed analysis of the full model is presented in Ref. supp .
Inferring hidden quantities. More often than not, experiments on small nonequilibrium systems lack the spatial or temporal resolution to monitor all relevant degrees of freedom, rendering the measurement of thermodynamic quantities very difficult. For this reason the study of counting observables through hidden subsets of states has received a lot of attention rahavFluctuationRelationsCoarsegraining2007 ; boEntropyProductionStochastic2014 ; biskerHierarchicalBoundsEntropy2017 ; yuDissipationLimitedResolutions2024a ; espositoStochasticThermodynamicsCoarse2012 ; ehrichTightestBoundHidden2021 ; blomMilestoningEstimatorsDissipation2024 ; liangThermodynamicBoundsSymmetry2024 ; Tan2021 , particularly entropy production. When a trajectory delves into hidden states, it performs an excursion until it resurfaces at one observable state, see Fig. 1(d); hence, harnessing properties of excursions aids the inference of properties such as violations of detailed balance, topology, and bounds on entropy production. Thus, a general framework for excursion observables shows promise to find applications in many inference problems. One particular example is inferring the energy displaced during an excursion from its duration.
Cell sensing. Membrane-bound receptors are fundamental tools that cells use to sense the concentration of chemicals in the environment and perform chemotaxis. Cues from these receptors are leveraged to climb food gradients, guide immune responses, and regulate gene expression. Errors in the inferred concentration were shown to satisfy lower bounds with some fundamental properties (such as the number of binding events) in cornerstone contributions bergPhysicsChemoreception1977 ; bialekPhysicalLimitsBiochemical2005 ; Mora2010 . More recently, a bound involving the energy expenditure was established Harvey2023 . In the latter, the estimation error was obtained from fluctuations in the time spent on signaling states, which is analogous to the duration of excursions through signaling states, see Fig. 1(e). Additionally, studying counting observables over these excursions, such as entropy production, can help to understand the energetic balances and the thermodynamic cost of the process. As an example application of our results, we explore the fluctuations of entropy production in a cell sensing model and a sub/super-Poissonian transition of excursion duration as functions of accuracy in Ref. supp .
Other scenarios. Excursions can also be identified as carriers of relevant observables in gene expression bursts wanDynamicImagingNascent2021 , cycle-resolved quantities of thermal machines Hegde2025 , periods of inactivity in queueing theory kleinrock1974queueing ; Lindley1952 ; gross2011fundamentals , search strategies tejedorOptimizingPersistentRandom2012 , the NP-hard problem of the traveling salesman, dynamics between metastable states, the history of a species before extinction, and the trajectory of a photon between successive scattering events.
Formalism—We consider a system that occupies a discrete alphabet of states and evolves stochastically according to a Markovian classical master equation. The probability $p_{x}$ of finding the system in state $x$ obeys
$\frac{dp_{x}}{dt}=\sum_{\ell}\sum_{y\neq x}W_{\ell xy}p_{y}-\Gamma_{x}p_{x},\qquad\Gamma_{x}=\sum_{\ell}\sum_{y\neq x}W_{\ell yx},$ (1)
where $\ell$ labels the different transitions between a pair of states  [^1] When different mechanisms induce the same change of states, they can be individually modeled by distinct transitions $\ell$, rendering the state-space a multigraph, and they can represent e.g. different reservoirs or chemical reactions., and $W_{\ell xy}$ is the transition rate to go from $y\to x$ through $\ell$. In vector notation, Eq. (1) can be written as $d|p\rangle/dt=\mathbb{W}|p\rangle$, where $|p\rangle$ is a vector with entries $p_{x}$ and
$\mathbb{W}=W-\Gamma=\begin{cases}W_{xy}&x\neq y\\ -\Gamma_{x}&x=y.\end{cases}$ (2)
Here and throughout $W$ is the matrix with entries $W_{xy}=\sum_{\ell}W_{\ell xy}$ for $x\neq y$, and zeros in the diagonal; conversely, $\Gamma$ is the diagonal matrix with $\Gamma_{x}$ in the diagonals. Assuming irreducibility of the state space, it follows that Eq. (1) has a unique steady state $|p^{\rm ss}\rangle$ which is the solution of $\mathbb{W}|p^{\rm ss}\rangle=0$. Eq. (1) describes the dynamics at the level of the ensemble. Conversely, we can also describe the process in a single stochastic trajectory such as $x_{1}\to_{\ell_{1}}x_{2}\to_{\ell_{2}}x_{3}\to_{\ell_{3}}\cdots$, where the system spends a random amount of time in a given state before jumping on to the next one. Such a trajectory is comprised of three elements: (i) a set of states $\{x_{1},x_{2},x_{3},\ldots\}$ that the system navigates; (ii) the residence times $\tau_{i}$, representing how much time the system spent in state $x_{i}$ before jumping to $x_{i+1}$; and (iii) the set of specific transitions $\ell_{i}$ that generated $x_{i}\to_{\ell_{i}}x_{i+1}$.
Suppose we split the alphabet of states into two arbitrary regions, denoted by $A$ and $B$. One may think of $A$ as some “inactive” phase and $B$ as an “active” one, but the splitting itself is absolutely general. We then define an excursion as a stochastic trajectory that begins whenever the system jumps from a state in $A\to B$ and ends whenever the system first returns from $B\to A$. In between, the system spends a random amount of time in $B$, and navigate over a random number of states in it, see Fig. 1(a). We emphasize that while Brownian excursions (i.e. random walks with fixed initial and final points) have been discussed in the literature before Chung1976 ; Majumdar2005 ; Majumdar2008 ; Majumdar2015 ; Takacs1991 ; Mazzolo2017 , our approach offers a generalization.
Counting observables—Our main goal in this Letter is to describe the statistics of counting observables within a single excursion. For one such single excursion, define a random variable $\hat{N}_{\ell xy}$ that counts how many times the transition $y\to_{\ell}x$ was observed (with $x,y\in A\cup B$). A counting observable is defined as a linear combination
$\hat{Q}=\sum_{\ell}\sum_{x,y}\nu_{\ell xy}\hat{N}_{\ell xy},$ (3)
with generic weights $\nu_{\ell xy}$ landi2024a . In FCS, counting observables are defined in a similar way, but they count the transitions over a fixed time interval $[0,t]$. Here, $\hat{Q}$ counts events during a single excursion, whose duration is also random. The counting observable (3) may count both transitions within $B$, as well as those between $A\leftrightarrow B$.
Choosing $\nu_{\ell xy}=1\,\forall\ell,x,y$ leads to a counting observable $\hat{\mathcal{A}}=\sum_{\ell,x,y}\hat{N}_{\ell xy}$ that addresses the total number of transitions within an excursion, which is the dynamical activity (freneticity) Di_Terlizzi2018 ; Nishiyama2024 ; Vo2022 . Similarly, $\nu_{\ell xy}=1\,\forall x,y$ for some $\ell$, can address the number of transitions mediated by e.g. a specific given reservoir or a chemical reaction.
Linear counting observables also describe thermodynamic currents. In this case the corresponding weights must be anti-symmetric $\nu_{\ell xy}=-\nu_{\ell yx}$ Wachtel2015 . Transition rates generated by thermal reservoirs satisfy local detailed balance: $W_{\ell xy}/W_{\ell yx}=\exp\{-\beta_{\ell}(E_{x}-E_{y})\},$ where $E_{x/y}$ are the energies of levels $x/y$ and $\beta_{\ell}$ is the inverse temperature of the reservoir associated with $\ell$. The counting observable for the entropy produced within a single excursion is $\hat{\Sigma}:=-\sum_{\ell,x,y}\beta_{\ell}(E_{x}-E_{y})\hat{N}_{\ell xy}$, corresponding to $\nu_{\ell xy}=-\beta_{\ell}(E_{x}-E_{y})$. Other thermodynamic currents, such as heat and work, can be constructed similarly.
Excursion statistics— Split the transition matrix (2) into a block structure
$\mathbb{W}=\begin{pmatrix}\mathbb{W}_{\!\!A}&W_{\!\!AB}\\ W_{\!\!BA}&\mathbb{W}_{\!\!B}\end{pmatrix},$ (4)
with the diagonal blocks $\mathbb{W}_{\!A}=W_{A}-\Gamma_{A}$ and $\mathbb{W}_{\!B}=W_{B}-\Gamma_{B}$. Note that $\mathbb{W}_{\!A}$ and $\mathbb{W}_{\!B}$ are not stochastic matrices since their columns do not add up to zero. Let $\hat{\mathcal{T}}$ denote the random duration of an excursion. Its PDF follows from $P(\hat{\mathcal{T}}=t)=E[\delta(t-\sum_{j=1}^{N}\tau_{j})]$, which results in
$P\big{(}\hat{\mathcal{T}}=t\big{)}=C_{x_{A}\to y_{A}}\left\langle y_{A}\left\lvert W_{AB}e^{\mathbb{W}_{\!B}t}W_{BA}\right\rvert x_{A}\right\rangle,$ (5)
where $C_{x_{A}\to y_{A}}^{-1}=\langle y_{A}|W_{AB}(-\mathbb{W}_{B}^{-1})W_{BA}|x_{A}\rangle$ is a normalization constant, which is proportional to the probability that an excursion follows the path $x_{A}\to y_{A}$ (see End Matter). Equation (5) is a minor generalization of known results on first passage times skinner2021 and a self-contained proof is presented in supp .
Next, consider a set of $r$ counting observables $\hat{Q}_{\alpha}$, defined as in Eq. (3), each with its own set of weights $\nu_{\ell xy}^{\alpha}$. Introduce tilted matrices
$(\mathbb{W}_{\bm{\xi}})_{xy}=\begin{cases}\sum_{\ell}W_{\ell xy}e^{i\sum_{\alpha}\nu_{\ell xy}^{\alpha}\xi_{\alpha}}&x\neq y\\[5.69046pt] -\Gamma_{x}&x=y,\end{cases}$ (6)
with $r$ counting fields $\xi_{\alpha}$. For an excursion starting in $x_{A}$ and ending in $y_{A}$, the joint probability that each observable $\hat{Q}_{\alpha}$ takes on a value $q_{\alpha}$ and that total excursion time is $\hat{\mathcal{T}}=t$ reads
$P(\bm{q},t)=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi_{1}\ldots d\xi_{r}}{(2\pi)^{r}}\langle y_{A}|W_{AB\bm{\xi}}e^{\mathbb{W}_{\!B\bm{\xi}}t}W_{BA\bm{\xi}}|x_{A}\rangle e^{-i\bm{q}\cdot\bm{\xi}},$ (7)
where $\bm{q}=(q_{1},\ldots,q_{r})$. All other calculations follow from this result. The proof is in Ref. supp with the rationale as follows. The counting fields $\bm{\xi}=(\xi_{1},\ldots,\xi_{r})$ pick up all events that occur during an excursion. We first count the starting jump from $A\to B$ using $W_{BA\bm{\xi}}$. Then we count an arbitrary number of jumps within $B$ using $\exp\{\mathbb{W}_{\!B\bm{\xi}}t\}$. And finally we count the jump from $B\to A$ with $W_{AB\bm{\xi}}$. With the tilted transition matrix (6), each transition $y\to x$ (via $\ell$) picks up the correct factor of $\nu_{\ell xy}^{\alpha}$ for each counting field. The integral over all $\xi_{\alpha}$ in Eq. (7) then takes us from the characteristic function to the actual probability distribution. Marginalizing Eq. (7) over $\bm{q}$ reduces back to Eq. (5).
Simple formulas for analytically calculating the moments of $\hat{Q}_{\alpha}$ and $\hat{\mathcal{T}}$, as well as their correlations, are provided in Ref. supp . We also provide details on how to evaluate some counting observables of interest such as the dynamical activity. Eq. (7) considers excursions between two specific states |$\ket{x_{A}}\to\ket{y_{A}}$ in $A$. To analyze the steady state behavior one needs simply to replace |$\ket{x_{A}}\to\ket{p^{\textrm{ss}}_{A}}$ and $\bra{y_{A}}=\bra{1_{A}}$ |, where |$\ket{p^{\textrm{ss}}_{A}}$ is the part of |$\ket{p^{\rm ss}}$ pertaining to region $A$ and $\bra{1_{A}}$ | is a vector with the dimensions of $A$, with all entries equal to 1. This therefore amounts to averaging the initial state over all possibilities, and marginalizing over all possible final states.
As a side note, we mention that the entropy production $\hat{\Sigma}$ satisfies an exchange fluctuation theorem at the level of individual excursions Jarzynski1997 ; Jarzynski2004 ; seifertEntropyProductionStochastic2005a ; DeRoeck2007 . Let $P_{x_{A}\to y_{A}}(\sigma)$ denote Eq. (7) for the probability of $\hat{Q}=\hat{\Sigma}$ assuming $\sigma$, and marginalized over $t$. We show in Ref. supp that, for a single excursion,
$\frac{P_{x_{A}\to y_{A}}(\sigma)}{P_{y_{A}\to x_{A}}(-\sigma)}=\frac{C_{x_{A}\to y_{A}}}{C_{y_{A}\to x_{A}}}e^{\sigma}.$ (8)
The factor of $C_{x_{A}\to y_{A}}/C_{y_{A}\to x_{A}}$ represents the ratio between the probability of an excursion $y_{A}\to x_{A}$ starting from $y_{A}$ and its opposite, and appears because the probabilities on the left-hand side are conditional on both excursion ends. For excursions that start and end in the same state $x_{A}=y_{A}$, or if we consider excursions regardless of start and end, this reduces to the usual exchange fluctuation theorem.
Connection with steady state FCS—We now connect the statistics of single excursions with steady state FCS. Let $\hat{\mathcal{Q}}$ denote a counting observable for the latter. It is defined similarly to Eq. (3), but integrated over a fixed time window $[0,t]$ instead of over a single excursion. In the limit of large $t$, one then defines the average current and the diffusion coefficient (noise)
$J=\lim\limits_{t\to\infty}\frac{E(\hat{\mathcal{Q}})}{t},\qquad D=\lim\limits_{t\to\infty}\frac{{\rm var}(\hat{\mathcal{Q}})}{t}.$ (9)
To establish the connection with excursions, we assume that region $A$ is composed of a single state, rendering excursions statistically independent of each other. Consider the concatenation of several excursions, one after the other, all starting in the same state $x_{A}$. Let $\hat{Q}_{n},\hat{\mathcal{T}}_{n}$ denote a counting observable and the excursion time during the $n$-th excursion. The total cycle time of an excursion is $\hat{\mathcal{T}}_{n}^{\rm cyc}=\hat{\mathcal{T}}_{n}+\hat{\tau}_{x,n}$, and should include the residence time $\hat{\tau}_{x,n}$ that the system spends in $x$ between excursions $n-1$ and $n$. Note that $\hat{\tau}_{x,n}$ and $\hat{\mathcal{T}}_{n}$ are statistically independent. Finally, let $\hat{\mathcal{N}}(t)$ denote the number of excursions that took place in the interval $[0,t]$. This random variable forms a renewal process cox1967renewal . For sufficiently large $t$, we can then write $\hat{\mathcal{Q}}\simeq\sum_{n=1}^{\hat{\mathcal{N}}(t)}\hat{Q}_{n}$, with the global constraint $\sum_{n=1}^{\hat{\mathcal{N}}(t)}\hat{\mathcal{T}}_{n}^{\rm cyc}\simeq t$. The error is manifested only in the boundary terms (e.g. at time $t$ it may be that $N$ excursions took place, but the system is still halfway through the ($N+1$)-th excursion). It is therefore sub-extensive and can be discarded in the long $t$ limit.
With this construction in place, we now state the main results (refer to the End Matter for the proofs). Let $\mu=E(\hat{\mathcal{T}}_{n}^{\rm cyc})=E(\hat{\mathcal{T}})+\Gamma_{x}^{-1}$ and $\Delta^{2}={\rm var}(\hat{\mathcal{T}}_{n}^{\rm cyc})={\rm var}(\hat{\mathcal{T}})+\Gamma_{x}^{-2}$ denote the mean and variance of the cycle time, with $E(\hat{\mathcal{T}})$ and ${\rm var}(\hat{\mathcal{T}})$ computed from Eq. (5). Here we also used the fact that $\hat{\tau}_{x}$ is exponentially distributed with parameter $\Gamma_{x}$. The steady state current in (9) is given by
$J=\frac{E(\hat{Q})}{\mu},$ (10)
with $E(\hat{Q})$ computed from Eq. (7). Next, and much less intuitively, the diffusion coefficient in (9) is given by
$D=\frac{{\rm var}(\hat{Q})}{\mu}+\frac{\Delta^{2}}{\mu^{3}}E(\hat{Q})^{2}-\frac{2E(\hat{Q})}{\mu^{2}}{\rm cov}(\hat{Q},\hat{\mathcal{T}}).$ (11)
It provides a decomposition of the FCS noise in terms of statistical contributions of individual excursions. The first two terms are what one might expect from the law of total variance, together with some known results from renewal theory cox1967renewal . The third term, however, is a correction term that appears due to the correlations between the counting observable and the excursion duration. This illustrates one of the triumphs of the framework presented here. By separately studying the three components of Eq. (11) one can shed light on the mechanisms responsible for the noise in FCS.
Application: Absorption refrigerator—To illustrate our framework, we apply it to characterize distributions and fluctuations of a three-qubit absorption refrigerator. Here we consider a simplification with only five energy states. 00 (resp. $1$) represents that the corresponding qubit is in the ground state (resp. is excited), see Fig. 1(b). Region $A$ is taken to be spanned only by the state $000$. The quantum dynamics is very well approximated supp by the classical master equation (1) in the parameter regime of Ref. aamirThermallyDrivenQuantum2025 .
Figure 2: (a) Covariance between cold current and excursion duration as a function of $n_{c}$, for different values of $\Gamma_{c}$ with fixed $g=10$. (b) Probability distribution in log scale of the dynamical activity. We fixed $g=10$ and $n_{c}=0.05$. (c) Waiting time distributions as a function of time for the first transition within $B$, excluding returning transitions (which would end the excursion). We fixed $g=10$ and $n_{c}=0.05$. (d), (e) Diffusion coefficient and its decomposition for (d) entropy production, and (e) dynamical activity as a function of $g$, with $n_{c}=0.1$ and $\Gamma_{c}=0.1$. For all plots, we fixed $\Gamma_{h}=50,\ \Gamma_{w}=0.5,\ n_{h}=0.005,\ n_{w}=0.5$. Parameters are well in accordance with the parameter regime of aamirThermallyDrivenQuantum2025 and the classical approximation.
Although there are many interesting counting observables to be explored in this model, we restrict ourselves to the cold current, entropy production, and dynamical activity. The cold current $\hat{J}_{c}$ is obtained by setting weights $\nu_{\ell xy}=\pm 1$ in Eq. (3) if the transition $x\to_{\ell}y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. Excursions can fully characterize the probability distribution of trajectories having zero cold current, in contrast with the steady state analysis supp . In Fig. 2(a) we show the covariance between the cold current $\hat{J}_{c}$ and the excursion duration $\hat{\mathcal{T}}$ as a function of the cold bath occupation $n_{c}$. Interestingly, longer excursions tend to have larger $\hat{J}_{c}$ in the cooling window, which is not necessarily the case when $\langle\hat{J}_{c}\rangle>0$, and a vanishing covariance pinpoints the value of $n_{c}$ where the cooling window begins, see End Matter. In Fig. 2(b) we show the distribution of the dynamical activity $\hat{\mathcal{A}}$, revealing that shorter excursions dominate the dynamics. Note that dynamical activity is lower bounded by $2$ (which is the shortest excursion length), and they are always even due to the specific model topology and allowed transitions in $B$. We go a step further and show in Fig. 2(c) the waiting-time distributions of the first transition within $B$ provided it does not return to $A$. Interestingly, the transition associated with the three-body coupling ($10\to 10$ 101) is dominant for shorter times, but is strongly suppressed afterward.
In Figs. 2(d) and (e), we characterize the diffusion coefficient (11) in terms of its components for the entropy production $D_{\Sigma}$ and the dynamical activity $D_{\mathcal{A}}$. In the former, the fluctuations are strongly dominated by the variance term of $\hat{\Sigma}$; whereas in the latter there is a clear competition between different contributions, with $E(\hat{\mathcal{A}})^{2}$ being the most relevant term. This result sheds light on the very nature of the noise in FCS. $\hat{\Sigma}$ is an observable that can have positive and negative values related by fluctuation theorems, whereas $\hat{\mathcal{A}}$ is strictly positive. The stark contrast of behavior on the two observables shows the versatility of the excursion framework to characterize fluctuations.
Conclusions—We have discussed a framework for stochastic excursions to investigate sub-trajectory statistics in systems far from equilibrium. By focusing on statistics of counting observables—such as currents, entropy production, and dynamical activity—during individual excursions, we are able to capture their fluctuations behavior at the sub-trajectory level. In particular, entropy production satisfies an exchange fluctuation theorem. Our central result, the decomposition of the diffusion coefficient beyond the law of total variance, reveals a fundamental connection between steady state noise described by FCS and fluctuations at the single excursion level. These results offer a new approach to analyze noise and irreversibility in classical stochastic systems.
There are many questions to be addressed as a follow-up to the work presented here. Initially, we point out that the excursions themselves may be correlated when region $A$ has more than one state, and such correlations may be a useful resource. We considered classical dynamics, but the concept of excursions itself is general and could be used in quantum settings as well. Finally, exploring the role of excursions in generic fluctuation theorems and uncertainty relations, as well as inference of counting observables, is subject to further work.
Acknowledgements.
GF acknowledges useful discussions with Felipe Pereira-Alves, Felipe D. Picoli and Jiheng Duan. This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. PH was supported by the project INTER/FNRS/20/15074473 funded by F.R.S.-FNRS (Belgium) and FNR (Luxembourg).